OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 12 in. and radius of 9 in.? Use 3.14 to approximate pi, and rexpress your final answer to the nearest tenth.

OpenStudy (anonymous):

@johnweldon1993 or anyone

OpenStudy (johnweldon1993):

So the equation would be \[\large V = \frac{\pi r^2 h}{3}\] So plugging everything in we have \[\large V = \frac{3.14\times 9^2 \times 12}{3}\] What does that come out to equal?

OpenStudy (anonymous):

35532.2

OpenStudy (johnweldon1993):

Not quite...so we have \[\large \frac{3.14 \times 81 \times 12}{3}\]

OpenStudy (anonymous):

31283.82 @johnweldon1993

OpenStudy (anonymous):

1017.36